OpenStudy (anonymous):

The owner of a gift shop purchases china figurines at a wholesale price of $3.80 each. The owner then displays each figurine in his shop with a sale price that is 75% higher than the wholesale price. After a period of time the shop owner decides to place the figurines on sale for 1/5 off their displayed price. To the nearest cent, what will be the new price for each figurine, after the discount?

OpenStudy (texaschic101):

purchased for 3.80 75% higher... 3.80 + .75(3.80) = 3.80 + 2.85 = 6.65 -- display price 1/5 off of display price 6.65 - 1/5(6.65) = 6.65 - 6.65/5 = 6.65 - 1.33 = 5.32 -- after discount price

OpenStudy (anonymous):

Thanks! Medal for you! :-)

OpenStudy (texaschic101):

:) happy,happy,happy